OpenStudy (anonymous):

Solve: u2 + 2u - 3 = 0

OpenStudy (anonymous):

Ok, to solve, you have to factor, do you know how to factor?

geerky42 (geerky42):

HINT: u²+2u-3 = (u + __)(u - __)

geerky42 (geerky42):

the sum of numbers in blank is 2 and product is -3. can you figure out what number they are?

OpenStudy (anonymous):

remember this formula (ax)^2 +(b+c)(ax) + bc =(ax+b)(ax+c)

geerky42 (geerky42):

@FallingBear